OpenStudy (horsegal244):

A bag contains 40 marbles, 4 of which are blue, 10 are red, 25 are green, and 1 is purple. Shawna takes a marble out of the bag, records the color, and returns it to the bag. How many green marbles should she expect after 400 trials?

OpenStudy (horsegal244):

@Michele_Laino @hartnn

OpenStudy (michele_laino):

the experimental frequency to get a green marble is: 25/40= 5/8 so the requested number is: \[\Large \frac{5}{8} \cdot 400 = ...?\]

OpenStudy (horsegal244):

640?

OpenStudy (michele_laino):

hint: \[\Large \frac{5}{8} \cdot 400 = \frac{{5 \cdot 400}}{8} = ...?\]

OpenStudy (horsegal244):

\[\frac{ 2000 }{ 8 }\]= 25

OpenStudy (michele_laino):

more precisely it is 250

OpenStudy (horsegal244):

ok

OpenStudy (horsegal244):

Can u help me with 3 more?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

here the experimental probability, is given by the subsequent ration: favorable outcomes/total outcomes: \[\Large \frac{{272}}{{600}} = ...?\]

OpenStudy (michele_laino):

ratio*

OpenStudy (horsegal244):

2.205

OpenStudy (michele_laino):

whereas the theoretical probability is: favorable outcomes (= 3, since there are three odd numbers) / possible outcomes(=6, since there are six possible outcomes): \[\Large \frac{3}{6} = ...?\]

OpenStudy (horsegal244):

2?

OpenStudy (horsegal244):

ohhh 1/2 duhhhh

OpenStudy (michele_laino):

we have: 272/600=0.453 and 3/6=0.5 so what can you conclude?

OpenStudy (horsegal244):

soit is not likely

OpenStudy (michele_laino):

we can see that experimental probability is less than the relative frequency

OpenStudy (horsegal244):

ok Next one ill start a new page...

OpenStudy (michele_laino):

ok!